Systems and Network Management

The MIB Tree

Figure 1 shows the Structure of Management Information (sim) object tree. It is also called a Management Information Base (mib) tree.

Figure 1: mib tree.

Notice that our Department has an official Enterprise number. This allows us to define our own object identifiers which can support both our own snmp mib tree, and our own ldap (Lightweight Directory Access Protocol) schemas. You can see our public listing

here: http://www.isi.edu/in-notes/iana/assignments/enterprise-numbers

Nick Urbanik <nicku@vtc.edu.hk>  ver. 1.0